A sign with a weight of 80N hangs from a vertical cable tied to 2 other cables. The upper cables make angles of 37 degrees and 53 degrees with the horizontal. Find the tension force in each cable.
Answer #1

Consider the tension in the first cable as T1 and the tension in the second cable as T2
Equating the horizontal forces,
T1 cos\theta1 = T2 cos\theta2
Where \theta 1 and \theta 2 are the angles 37 and 53 respectively.
T1 cos(37) = T2 cos(53)
T1 x 0.798 = T2 x 0.602
T1 = 0.754 T2          ...(1)

Equating the vertical forces,
T1 sin\theta1 + T2 sin\theta2 = 80
T1 x 0.612 + T2 x 0.798 = 80
Substituting T1,
0.754 T2 + 0.798 T2 = 80
T2 = 51.54 N

T1 = 0.754 T2
= 38.84 N